# Complexity of deep computations via topology of function spaces

### v2

@@ 2.5. Finite precision prefix test
Fix $a\in 2^{\mathbb{N}}$. Let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$. We claim that $\lim_{n\rightarrow\infty}\psi_{w_{n}}(x)=\delta_{a}(x)$ for all $x\in L$. If $x=a$, then $x|_{|w_{n}|}=w_{n}$ for all $n$ and so $\psi_{w_{n}}(x)=1^{\infty}=\delta_{a}(x)$ for all $n$. If $x\neq a$, then $x|_{|w_{n}|}\neq w_{n}$ for large enough $n$. Hence, $\psi_{w_{n}}(x)=0^{\infty}=\delta_{a}(x)$ for large enough $n$. ∎
@@
To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$ consider, for each $n\in\mathbb{N}$, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0. If $x=1^{\infty}$, then $x|_{|w_{n}|}\neq w_{n}$ for all $n\in\mathbb{N}$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for all $n\in\mathbb{N}$. If $x\neq 1^{\infty}$, let $N$ be the smallest such that $x(N)=0$. Then, $x|_{|w_{n}|}\neq w_{n}$ for all $n>N$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for large enough $n$. ∎
@@ 3.1. NIP, Rosenthal compacta, and deep computations
    $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$.
@@
3.2. The Todorčević trichotomy and levels of PAC learnability
@@
Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9). Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99], and later Argyros, Dodos, Kanellopoulos [ADK08] proved an heptachotomy that refined Todorčević’s classification. In this section, inspired by the work of Glasner and Megrelishvili [GM22], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
@@
We now present some separable and non-separable examples of Rosenthal compacta. These show that the previously discussed classes NIPi are not equal.
@@
    Alexandroff compactification of a discrete space of size continuum. For each $a\in 2^{\mathbb{N}}$ consider the map $\delta_{a}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ given by $\delta_{a}(x)=1$ if $x=a$ and $\delta_{a}(x)=0$ otherwise. Let $A(2^{\mathbb{N}})=\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{0\}$, where 0 is the zero map. Notice that $A(2^{\mathbb{N}})$ is a compact subset of $B_{1}(2^{\mathbb{N}})$, in fact $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$ is a discrete subspace of $B_{1}(2^{\mathbb{N}})$ and its pointwise closure is precisely $A(2^{\mathbb{N}})$. Hence, this is a Rosenthal compactum which is not first countable. Notice that this space is also not separable.
@@
    Split Cantor. Let << be the lexicographic order in the space of infinite binary sequences, i.e., $2^{\mathbb{N}}$. For each $a\in 2^{\mathbb{N}}$ let $f_{a}^{-}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{-}(x)=1$ if $x<a$ and $f_{a}^{-}(x)=0$ otherwise. Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise. The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$, which was obtained as the closure of the space discussed in Section 2.4, giving an example separating NIP2 from NIP3. This is a separable Rosenthal compactum. One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant. Moreover, it is hereditarily separable, but it is not metrizable.
@@
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a), if x∈C(K);0, if x∈K; g_a^1(x)={x(a), if x∈C(K);δa(x), if x∈K. Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable. The interesting case will be when $K=2^{\mathbb{N}}$.
@@
    Extended Alexandroff Duplicate of the split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x. Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace (see Theorem 5 in [TOD99]).
@@
    Let $(L,\mathcal{P},\Gamma)$ be the finite precision prefix test (Section 2.5) with $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$. We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3(3)), which is separable but not first countable. Hence, $\Delta$ is NIP but not NIP1.
@@
    Our emphasis is computational. Elements of $2^{<\mathbb{N}}$ represent finite bitstrings, i.e., standard computations, while Rosenthal compacta represent deep computations, i.e., limits of finite computations. Mathematically, deep computations are pointwise limits of standard computations. However, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
@@
    If $I$ is a countable and $\{f_{i}:i\in I\}\subseteq\mathbb{R}^{X}$, $\{g_{i}:i\in I\}\subseteq\mathbb{R}^{X}$ are two pointwise families by $I$, we say that $\{f_{i}:i\in\ I\}$ and $\{g_{i}:i\in I\}$ are equivalent if and only if the map $f_{i}\mapsto g_{i}$ is extended to a homeomorphism from $\overline{\{f_{i}:i\in I\}}$ to $\overline{\{g_{i}:i\in I\}}$.
@@
One of the main results in [ADK08] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families. We shall describe the seven minimal families next. We follow the same notation as in [ADK08]. For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s). Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$. Given $t\in 2^{<\mathbb{N}}$, let $v_{t}$ be the characteristic function of the set $\{x\in 2^{\mathbb{N}}:x\text{ extends }t\}$. Let << be the lexicographic order in $2^{\mathbb{N}}$. Given $a\in 2^{\mathbb{N}}$, let $f^{+}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a\leq x\}$ and let $f^{-}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a<x\}$. Given two maps $f,g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ we denote by $(f,g):2^{\mathbb{N}}\sqcup 2^{\mathbb{N}}\rightarrow\mathbb{R}$ the function which is $f$ on the first copy of $2^{\mathbb{N}}$ and $g$ on the second copy of $2^{\mathbb{N}}$.
@@
    $D_{3}=\{f^{+}_{s_{t}^{\frown}0^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is a discrete in $\overline{D_{3}}=S(2^{\mathbb{N}})$.
@@
    $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$
@@
The implications of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximate all the other deep computations. For example: in the finite precision prefix test example (subsection 2.5), the prefix test computations (family $D_{5}$) approximate all other deep computations. However, note that this discrete set $D_{i}$ may not consist of continuous functions (i.e., they will not be computable in general). For example, functions in $D_{3}$ are not continuous.

### v3

@@ 2.5. Finite precision prefix test
Fix $a\in 2^{\mathbb{N}}$, and let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$. We claim that $\lim_{n\rightarrow\infty}\psi_{w_{n}}(x)=\delta_{a}(x)$ for all $x\in L$. If $x=a$, then $x|_{|w_{n}|}=w_{n}$ for all $n$ and so $\psi_{w_{n}}(x)=1^{\infty}=\delta_{a}(x)$ for all $n$. If $x\neq a$, then $x|_{|w_{n}|}\neq w_{n}$ for large enough $n$. Hence, $\psi_{w_{n}}(x)=0^{\infty}=\delta_{a}(x)$ for large enough $n$. ∎
@@
To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$, for each $n\in\mathbb{N}$, consider, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0. If $x=1^{\infty}$, then $x|_{|w_{n}|}\neq w_{n}$ for all $n\in\mathbb{N}$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for all $n\in\mathbb{N}$. If $x\neq 1^{\infty}$, let $N$ be the smallest such that $x(N)=0$. Then, $x|_{|w_{n}|}\neq w_{n}$ for all $n>N$. Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for large enough $n$. ∎
@@ 3.1. NIP, Rosenthal compacta, and deep computations
    $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$;
@@
3.2. The Todorčević trichotomy, and levels of NIP and PAC learnability
@@
Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1), the set $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9). Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99] that was later refined through an heptachotomy proved by Argyros, Dodos, Kanellopoulos [ADK08]. In this section, inspired by the work of Glasner and Megrelishvili [GM22], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
@@
    Alexandroff compactification of a discrete space of size continuum. For each $a\in 2^{\mathbb{N}}$ consider the map $\delta_{a}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ given by $\delta_{a}(x)=1$ if $x=a$ and $\delta_{a}(x)=0$ otherwise. Let $A(2^{\mathbb{N}})=\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{\mathbf{0}\}$, where $\mathbf{0}$ is the zero map. Notice that $A(2^{\mathbb{N}})$ is a compact subset of $B_{1}(2^{\mathbb{N}})$. In fact, $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$ is an uncountable discrete subspace of $B_{1}(2^{\mathbb{N}})$, and its pointwise closure is precisely $A(2^{\mathbb{N}})$. Hence, this is a Rosenthal compactum which is not hereditarily separable (and therefore not first countable). In particular, this space is does not satisfy separability, but it can be made separable by adding a countable set as the next example shows.
@@
    Split Cantor. Let << be the lexicographic order in the space of infinite binary sequences, i.e., $2^{\mathbb{N}}$. For each $a\in 2^{\mathbb{N}}$ let $f_{a}^{-}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{-}(x)=1$ if $x<a$ and $f_{a}^{-}(x)=0$ otherwise. Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise. The Split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$, which was obtained as the closure of the space discussed in Section 2.4, giving an example separating NIP2 from NIP3. This is a well known separable Rosenthal compactum. One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant. Moreover, it is hereditarily separable, but it is not metrizable. It is homeomorphic to the space $2^{\mathbb{N}}\times\{0,1\}$ with the lexicographic order topology via the identification $(a,1)\mapsto f_{a}^{+}$ and $(a,0)\mapsto f_{a}^{-}$.
@@
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={0, if x∈Kx(a), if x∈C(K); g_a^1(x)={δa(x), if x∈K;x(a), if x∈C(K). Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $g_{a}^{i}:a\in U,i\in\{0,1\}\}\backslash\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set. Another abstract way in which this space is presented is as the space $K\times\{0,1\}$ whose basic open neighborhoods are given as before, identifying $(a,0)\mapsto g_{a}^{0}$ and $(a,1)\mapsto g_{a}^{1}$. We can also embed $D(K)$ into the product $A(K)\times K$ by identifying $(a,0)\mapsto(\mathbf{0},a)$ and $(a,1)\mapsto(\delta_{a},a)$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable, thus we typically study the interesting case when $K=2^{\mathbb{N}}$. As with the Alexandroff compactification $A(2^{\mathbb{N}})$, we can make the space $D(2^{\mathbb{N}})$ separable by adding a countable set. For example, the closure of the set $\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}\subseteq\hat{A}(2^{\mathbb{N}})\times 2^{\mathbb{N}}$ is $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}\cup\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}$, where $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}$ is homeomorphic to $D(2^{\mathbb{N}})$.
@@
    Extended Alexandroff Duplicate of the Split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x. Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. The identification $h_{t}\mapsto(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ lifts to a homeomorphism between $\hat{D}(S(2^{\mathbb{N}}))$ and the subspace of $\hat{A}(2^{\mathbb{N}})\times S(2^{\mathbb{N}})$ consisting of $(\mathbf{0},f_{a}^{+})$, $(\mathbf{0},f_{a}^{-})$, $(\delta_{a},f_{a}^{+})$ and $(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ for $a\in 2^{\mathbb{N}}$ and $t\in 2^{<\mathbb{N}}$ (see 4.3.7 in [ADK08]). Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace.
@@
    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by γ_t(x)={0, if lext0;(01), if t0lexx ≤lext1;1, if t1¡lexx. where $(01)^{\infty}$ denotes the sequence of alternating bits: $010101\cdots$. As in the other examples, it is not difficult to see that $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. For example, the condition $t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty}$ is equivalent to $x$ extending $t$. Observe that the set of deep computations is homeomorphic to $\hat{D}(S(2^{\mathbb{N}}))$ (see Example 3.3(5)). This is an example of $\Delta$ which is NIP1 but not NIP2.
  4. (4)
    Let $(L,\mathcal{P},\Gamma)$ be the finite precision prefix test (Section 2.5) with $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$. We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3-(3)), which is separable but not first countable. Hence, $\Delta$ is NIP but not NIP1.
@@
    Our emphasis is computational. Real numbers can be represented as infinite binary sequences, i.e., infinite branches of the binary tree $2^{<\mathbb{N}}$. We approximate real numbers or binary sequences with elements in $2^{<\mathbb{N}}$, i.e., finite bitstrings. Indexing standard computations with finite bitstrings allow us to better understand how deep computations arise and how they get approximated. Computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
@@
    If $I$ is countable and $\{f_{i}:i\in I\}\subseteq\mathbb{R}^{X}$, $\{g_{i}:i\in I\}\subseteq\mathbb{R}^{X}$ are two pointwise families by $I$, we say that $\{f_{i}:i\in\ I\}$ and $\{g_{i}:i\in I\}$ are equivalent if and only if the map $f_{i}\mapsto g_{i}$ is extended to a homeomorphism from $\overline{\{f_{i}:i\in I\}}$ to $\overline{\{g_{i}:i\in I\}}$.
@@
One of the main results in [ADK08] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families. We shall describe the seven minimal families next. We follow the same notation as in [ADK08]. For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing with all 0’s (respectively, all $1$’s). Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$. Given $t\in 2^{<\mathbb{N}}$, let $v_{t}$ be the characteristic function of the set $\{x\in 2^{\mathbb{N}}:x\text{ extends }t\}$. Let << be the lexicographic order in $2^{\mathbb{N}}$. Given $a\in 2^{\mathbb{N}}$, let $f^{+}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a\leq x\}$ and let $f^{-}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a<x\}$. Given two maps $f,g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ we denote by $(f,g):2^{\mathbb{N}}\sqcup 2^{\mathbb{N}}\rightarrow\mathbb{R}$ the function which is $f$ on the first copy of $2^{\mathbb{N}}$ and $g$ on the second copy of $2^{\mathbb{N}}$.
@@
    $D_{3}=\{f^{+}_{s_{t}^{\frown}0^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{3}}=S(2^{\mathbb{N}})$.
@@
    $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$.
@@
The implication of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximates all the other deep computations. For example: in the finite precision prefix test example (subsection 2.5), the prefix test computations (family $D_{5}$) approximate all other deep computations. However, note that this discrete set $D_{i}$ may not consist of continuous functions (i.e., they will not be computable in general). For example, functions in $D_{3}$ are not continuous.